9. Consider the three stocks in the following table. P, represents price at time t, and Q, represents shares outstanding at time t. Stock C splits two-for-one in the last period. (LO 2-2) Pt 95 45 110 100 200 200 100 200 400 95 90 50 100 100 200 200 45 a. Calculate the rate of return on a price-weighted index of the three stocks for the first period (i 0 to t 1). b. What must happen to the divisor for the price-weighted index in year 2? c. Calculate the rate of return of the price-weighted index for the second period (1o t = 2).
Answer #1

a.

P0 P1 P2
A 90 95 95
B 50 45 45
C 100 110 55
Average 80 83.3333 65

Price Weighted Return from t0 to t1 = (83.33333-80)/80 = 4.17%.

b. New Divisor for period 2.

P1 Spilt P2
A 95 95
B 45 45
C 110 2 55
Total 250 195

Average of P1 = 250/3 =83.333,

New Divisor = 195/83.33 = 2.34.